# Weyl metrizability of 3-dimensional projective structures and CR submanifolds

### v2

@@ 4.1. Projective class of a conformal Weyl structure
Here we obtain necessary conditions for the Weyl metrizability of an oriented projective structure.
@@
where the functions $\mathsf{P}_{ij}$ on $\mathcal{C}_{0}$ are referred to as the Schouten tensor or the Rho tensor of $\nabla^{[g]}.$ Straightforward computation shows, c.f. the proof of Theorem 5.2, that the pull-back of the Cartan connection $\varphi$ for the projective structure $[\nabla^{[g]}]$ to $\mathcal{C}_{0},$ in matrix form (3.6), is given by
@@
where, defining $[\varepsilon_{ij}]$ such that $\varepsilon_{00}=1,\varepsilon_{11}=\varepsilon_{22}=\varepsilon$ and $\varepsilon_{ij}=0$ for $i\neq j,$ one has
@@
Let $\mathbb{P}_{+}T^{*}M$ denote the bundle of rays of future time-like 1-forms in the Lorentzian signature and the oriented projectivized cotangent bundle in the Riemannian signature. The following holds.
@@
If an oriented projective structure []delimited-[][$\nabla$][ ∇ ] on a 3-manifold $M$ is Weyl metrizable for a conformal structure $[g]$ then there exists a section $s\colon\mathbb{P}_{+}T^{*}M\to\mathcal{T}^{[\nabla]}$ with the property that $\mathbb{P}_{+}T^{*}M$ is a CR submanifold of $\mathcal{T}^{[\nabla]}$ with an induced nondegenerate Levi bracket and of hypersurface type whose fundamental binary quartic is either zero or has a repeated root of multiplicity four. Moreover, in terms of the CR coframing (3.13), on $N$ one has the nonintegrability condition
@@
where $\varepsilon=1$ or $-1$ if $[g]$ is Riemannian or Lorentzian, respectively.
@@
Given a conformal connection $\nabla^{[g]}$ for $[g],$ by Proposition 3.2, the 2-nondegenerate CR structure on $\mathcal{T}^{[\nabla^{[g]}]}$ has holomorphic distribution ^$\widehat{\mathscr{H}}\subset\mathbb{C}\otimes T\mathcal{T}^{[\nabla^{[g]}]}$ given by $\mathrm{Ker}\{\tau^{0},\zeta^{1},\zeta^{2},\zeta^{3}\}$. Since the conformal connection $\nabla^{[g]}$ corresponds to a section $s\colon\mathcal{C}_{0}\to\mathcal{C}$, the projective connection $\nabla^{[g]}\in[\nabla^{[g]}]$ in turn defines an inclusion of principal bundles
@@
where $(\mathcal{P}\to M,\varphi)$ is the Cartan geometric data for the oriented projective structure $[\nabla^{[g]}].$
@@
(4.6) $\iota(\mathcal{T}^{[g]}):=\iota(\mathcal{C}_{0})/\mathrm{CO}(2,\mathbb{R})\subset\mathcal{T}^{[\nabla^{[g]}]}\cong\mathcal{P}/P_{J},$
wherein we are using the induced injection $\mathrm{CO}(2,\mathbb{R})\to P_{J};$ see § 3.3. Recalling that $\mathcal{T}^{[g]}\cong\mathbb{P}_{+}T^{*}M,$ (4.6) defines a section $\mathbb{P}_{+}T^{*}M\to\mathcal{T}^{[\nabla^{[g]}]}.$
The pull-back of $\varphi$ for $[\nabla^{[g]}]$ to $\mathbb{P}_{+}T^{*}M$ is expressed as (4.2). Using the definition of the contact 1-form $\tau^{0}$ and the holomorphic 1-forms $\zeta^{i}$ from (3.13), one obtains
@@
which, by (2.17), coincides with the contact and holomorphic 1-forms for the CR structure on $\mathcal{T}^{[g]}.$ In other words, $\iota(\mathcal{T}^{[g]})$ is a CR submanifold of $\mathcal{T}^{[\nabla^{[g]}]}$ that is transverse to its Levi kernel.
@@
A Weyl structure $\nabla^{[g]}$ for which $\mathsf{P}_{ij}=\mathsf{P}_{ji},$ is called a closed Weyl structure since, by structure equations (2.4), $\mathrm{d}\theta^{0}=0$ holds. Furthermore, given any section $s_{0}\colon M\to\mathcal{C}_{0},$ one has $\nabla^{[g]}s_{0}^{*}g=-2s_{0}^{*}\theta^{0}\otimes s_{0}^{*}g$ for $g\in\mathrm{Sym}^{2}(T^{*}\mathcal{C})$ in (2.2). Denoting $g_{0}=s_{0}^{*}g$ and $\phi_{0}=s^{*}_{0}\theta^{0},$ from $\mathrm{d}\phi_{0}=0,$ locally, $\phi_{0}=\mathrm{d}f$ holds for some function $f.$ It is then elementary to show that $\nabla^{[g]}g_{1}=0$ where $g_{1}=e^{2f}g_{0}$ and that $\nabla^{[g]}$ is the Levi-Civita connection of $g_{1}.$ In other words, closed Weyl structures are locally the Levi-Civita connection of representatives in $[g].$ The Rho tensor $\mathsf{P}_{ij}$ and the Ricci tensor ${R_{ij}}$ of the metric are related via $\mathsf{P}_{ij}=\tfrac{-1}{n-2}({R}_{ij}-\tfrac{1}{2(n-1)}R\varepsilon_{ij})$ where $R=\varepsilon^{ij}{R}_{ij}$ and $n=3.$
@@ 4.2. CR submanifolds as holomorphic sections of the Levi kernel
To show that the conditions in Proposition 4.1 are also sufficient, we need the following definition and an observation about Weyl structures for twistor CR manifolds wherein $\mathbb{P}_{+}T^{*}M$ denotes the bundle of rays of 1-forms.
@@
Given an oriented 3-dimensional projective structure []delimited-[][$\nabla$][ ∇ ] with twistor bundle $\mathcal{T}^{[\nabla]}\to\mathbb{P}_{+}T^{*}M,$ a section $\tilde{t}\colon N\to\mathcal{T}^{[\nabla]}$ for an open subset $N\subset\mathbb{P}_{+}T^{*}M$ is called holomorphic if the complex Pfaffian system $s^{*}\mathcal{I}=s^{*}\{\tau^{0},\zeta^{1},\zeta^{2},\zeta^{3}\},$ as defined in (3.13), is integrable.
By the construction in § 3.3, the fibers of $\mathcal{T}^{[\nabla]}\to\mathbb{P}_{+}T^{*}M$ are the leaves of the Pfaffian system $\{\tau^{0},\zeta^{1},\overline{\zeta}^{1},\zeta^{2},\overline{\zeta}^{2}\}.$ It follows that, for any section $\tilde{t}\colon N\to\mathcal{T}^{[\nabla]},$ $N\subset\mathbb{P}_{+}T^{*}M,$ the 1-forms $(s^{*}\tau^{0},s^{*}\zeta^{1},s^{*}\zeta^{2},s^{*}\overline{\zeta}^{1},s^{*}\overline{\zeta}^{2})$ give a coframe on $N$ and
$s^{*}\zeta^{3}=a_{0}\tau^{0}+a_{1}\zeta^{1}+a_{2}\zeta^{2}+b_{1}\overline{\zeta}^{1}+b_{2}\overline{\zeta}^{2},$
for some complex-valued functions $a_{0},a_{1},a_{2},b_{1},b_{2}$ on $N,$ where we have dropped $s^{*}$ on the right hand side. Since $\{\tau^{0},\zeta^{1},\zeta^{2},\zeta^{3}\}$ is integrable on $\mathbb{C}\otimes T^{*}\mathcal{T}^{[\nabla]},$ by Definition 4.3, $\tilde{t}\colon N\to\mathcal{T}^{[\nabla]}$ is holomorphic if and only if $b_{1}=b_{2}=0.$
@@
Recall from § 3.2 that an oriented projective structure, [],delimited-[][$\nabla$],[ ∇ ] , defines an almost para-CR structure on $\mathbb{P}_{+}T^{*}M,$ which is a Cartan geometry $(\mathcal{P}\to\mathbb{P}_{+}T^{*}M,\varphi)$ of type $(\mathrm{SL}(4,\mathbb{R}),P_{1,3}),$ where $P_{1,3}=P_{0}\ltimes P_{+}$ and $P_{0}=\mathbb{R}^{+}\times\mathrm{GL}^{+}(2,\mathbb{R}).$ Since by (3.11) the fibers of the twistor bundle $\mathcal{T}^{[\nabla]}\to\mathbb{P}_{+}T^{*}M$ are $\mathrm{SL}(2,\mathbb{R})/\mathrm{SO}(2,\mathbb{R})\cong\mathbb{D}^{2},$ a section $\tilde{t}\colon N\to\mathcal{T}^{[\nabla]},N\subset\mathbb{P}_{+}T^{*}M,$ is equivalent to a reduction of the structure group of the almost para-CR structure on $N$ from $P_{0}=\mathbb{R}^{+}\times\mathrm{GL}^{+}(2,\mathbb{R})$ to $\widetilde{P}_{0}:=\mathbb{R}^{+}\times\mathrm{CO}(2,\mathbb{R}).$ As a result of such reduction, one obtains an inclusion $\iota_{0}\colon\widetilde{\mathcal{P}}_{0}\to\mathcal{P}$ where $\widetilde{\mathcal{P}}_{0}$ is a principal $(\widetilde{P}_{0}\ltimes P_{+})$-bundle. Moreover, by Definition 4.3 and the subsequent discussion, since $\tilde{t}\colon N\to\mathcal{T}^{[\nabla]}$ is holomorphic, it follows that
$\tilde{t}^{*}\zeta^{3}=a_{0}\tau^{0}+a_{1}\zeta^{1}+a_{2}\zeta^{2},$
for some $a_{0},a_{1},a_{2}\in C^{\infty}(\widetilde{\mathcal{P}}_{0},\mathbb{C}).$ One can explicitly find the dependency of $a_{i}$’s on $\widetilde{P}_{0}\ltimes P_{+}$ as follows. If two points $p,q\in\widetilde{\mathcal{P}}_{0}$ lie on the fiber over $x\in N\subset\mathbb{P}_{+}T^{*}M,$ then there is $\mathbf{g}\in\widetilde{P}_{0}\ltimes P_{+}$ such that $q=\mathbf{g}^{-1}p,$ where
@@
$\varphi(p)\to\varphi(\mathbf{g}^{-1}p)=\mathbf{g}^{-1}\varphi\mathbf{g}+\mathbf{g}^{-1}\mathrm{d}\mathbf{g}.$
It is straightforward to find the transformations $\tau^{0}(p)\to\tau^{0}(\mathbf{g}^{-1}p)$ and $\zeta^{i}(p)\to\zeta^{i}(\mathbf{g}^{-1}p)$ to be
(4.9) $\begin{gathered}\tau^{0}\to\tfrac{1}{\mathbf{b}_{0}^{2}\mathbf{b}_{1}^{2}}\tau^{0},\quad\zeta^{1}\to\tfrac{e^{\mathrm{i}\mathbf{c}}}{\mathbf{b}_{0}^{3}\mathbf{b}_{1}}\zeta^{1}-\tfrac{\mathbf{q}_{34}}{\mathbf{b}_{0}^{2}\mathbf{b}_{1}^{2}}\tau^{0},\quad\zeta^{2}\to\tfrac{\mathbf{b}_{0}e^{\mathrm{i}\mathbf{c}}}{\mathbf{b}_{1}}\zeta^{2}+\tfrac{\mathbf{q}_{12}}{\mathbf{b}_{0}^{2}\mathbf{b}_{1}^{2}}\tau^{0},\\ \zeta^{3}\to e^{2\mathrm{i}\mathbf{c}}\zeta^{3}+\tfrac{\mathbf{b}_{0}e^{\mathrm{i}\mathbf{c}}}{\mathbf{b}_{1}}\zeta^{2}+\tfrac{e^{\mathrm{i}\mathbf{c}}}{\mathbf{b}_{0}^{3}\mathbf{b}_{1}}\zeta^{1}-\tfrac{\mathbf{q}_{12}\mathbf{q}_{34}}{\mathbf{b}_{0}^{2}\mathbf{b}_{1}^{2}}\tau^{0},\\ \end{gathered}$
@@
$\iota^{*}_{01}\zeta^{3}=a\tau^{0},$
for some $a\in C^{\infty}(\mathcal{P}_{1},\mathbb{C})$ where $\iota_{01}=\iota_{0}\circ\iota_{1}\colon\mathcal{P}_{1}\to\mathcal{P}.$ Setting $a_{1}=a_{2}=0$ in differential relations (4.10), one arrives at
$\iota_{01}^{*}\mu_{12}=$ $X_{0}\tau^{0}+X_{1}\zeta^{1}-Y_{1}\zeta^{2}-\textstyle{\frac{1}{2}}a_{0}\overline{\zeta}^{2}+\textstyle{\frac{1}{2}}(W^{1}_{212}+W^{2}_{121}+\mathrm{i}W^{1}_{112}+\mathrm{i}W^{2}_{221})\overline{\zeta}^{1},$
@@
for some complex-valued functions $X_{0},X_{1},Y_{0},Y_{1},Y_{2}$ on $\mathcal{P}_{1}.$
@@
Thus, one obtains a sub-bundle $\iota_{3}\colon\mathcal{P}_{3}\to\mathcal{P}_{2},$ which is a principal $\mathrm{CO}(2,\mathbb{R})\times\mathbb{R}^{+}$-bundle, defined as
(4.11) $\mathcal{P}_{3}=\left\{u\in\mathcal{P}_{2}\ \vline\ \Re Y_{1}(p)=0\right\}.$
Any holomorphic section $\tilde{t}\colon N\to\mathcal{T}^{[\nabla]},N\subset\mathbb{P}_{+}T^{*}M,$ defines a nondegenerate CR structure of hypersurface type on $N$ since $\mathscr{F}:=\mathrm{Ker}\{\tau^{0}\}$ is the canonical contact distribution and $\mathscr{H}:=\mathrm{Ker}\{\tilde{t}^{*}\mathcal{I}\}\subset\mathbb{C}\otimes\mathscr{F}$ is an integrable holomorphic distribution of a complex structure which by (3.15) satisfies the compatibility condition. In terms of the Cartan geometric description of Theorem 2.5, in the induced CR structure on $N\subset\mathbb{P}_{+}T^{*}M,$ the principal $\widetilde{P}_{0}$-bundle $\mathcal{P}_{3}\to N$ defines a CR (Weyl) connection $t\colon\mathcal{Q}_{0}\to\mathcal{Q}$ on $N$. Setting $\iota=\iota_{1}\circ\iota_{2}\circ\iota_{3}\colon\mathcal{P}_{3}\to\mathcal{P},$ one can express the entries of $\iota^{*}\gamma$ in (2.8) as
@@
The expression for $C_{4}$ in (4.12) is given in terms of the first jet the fundamental torsion (3.7) for the almost para-CR structure on $\mathbb{P}_{+}T^{*}M$ defined by [],delimited-[][$\nabla$],[ ∇ ] , which leads to Corollary 5.1. Furthermore, it follows from (4.9) that the property of being holomorphic in Definition 4.3 for a section $\tilde{t}\colon N\to\mathcal{T}^{[\nabla]},N\subset\mathbb{P}_{+}T^{*}M$ is invariant under the action of the structure group. Lastly, it is clear from $C_{0}=0$ in (4.12) that the binary quartic (2.12) has a real root at $[a_{1}:a_{2}]=[1:0]$ which, by the parametrization (2.11) and Proposition 2.9, corresponds to the fibers of $\mathbb{P}_{+}T^{*}M\to M.$

### v3

@@ 4.1. Projective class of a conformal Weyl structure
Here we obtain necessary conditions for the Weyl metrizability of an oriented projective structure. The Cartan geometry of a conformal structure $[g],$ is denoted by $(\mathcal{C}\to M,\psi),$ satisfying (2.1) and (2.3). The Cartan geometry of a projective structure, $\nabla$], is denoted as $(\mathcal{P}\to M,\varphi),$ satisfying (3.2).
@@
where the functions $\mathsf{P}_{ij}$ on $\mathcal{C}_{0}$ are referred to as the Schouten tensor or the Rho tensor of $\nabla^{[g]}.$ Subsequently, the inclusion $\mathcal{C}_{0}\subset\mathcal{P}$ gives a $\mathrm{CO}(p,q)$ reduction of the projective structure of $\nabla^{[g]}]$, i.e. a reduction of the Cartan bundle $\mathcal{P}\to M,$ to the principal $\mathrm{CO}(p,q)$-bundle $\mathcal{C}_{0}\to M$ defined by the conformal connection $\nabla^{[g]}.$ Straightforward computation shows, c.f. the proof of Theorem 5.2, that the pull-back of the Cartan connection $\varphi$ for the projective structure $\nabla^{[g]}]$ to $\mathcal{C}_{0},$ in matrix form (3.6), is given by
@@
where, defining $\varepsilon_{ij}]$ such that $\varepsilon_{00}=1,\varepsilon_{11}=\varepsilon_{22}=\varepsilon$ and $\varepsilon_{ij}=0$ for $i\neq j,$ one has
@@
Let $\mathcal{R}^{\uparrow}T^{*}M$ denote the bundle of rays of future time-like 1-forms in the Lorentzian signature and the oriented projectivized cotangent bundle in the Riemannian signature. The following holds.
@@
If an oriented projective structure $\nabla$] on a 3-manifold $M$ is Weyl metrizable for a conformal structure $[g]$ then there exists a section $s\colon\mathcal{R}^{\uparrow}T^{*}M\to\mathcal{T}^{[\nabla]}$ with the property that $s(\mathcal{R}^{\uparrow}T^{*}M)$ is a CR submanifold of $\mathcal{T}^{[\nabla]}$ with an induced nondegenerate Levi bracket and of hypersurface type whose fundamental binary quartic is either zero or has a repeated root of multiplicity four. Moreover, in terms of the CR coframing (3.13), on $N$ one has the nonintegrability condition
@@
where $\varepsilon=1$ or 1-1 if $[g]$ is Riemannian or Lorentzian, respectively.
In the statement of the proposition, since $\mathcal{R}^{\uparrow}T^{*}M\subseteq\mathbb{P}_{+}T^{*}M$ is an open subset, by restriction, we have the fibration $\mathcal{T}^{[\nabla]}\to\mathcal{R}^{\uparrow}T^{*}M$ using which we can define sections $s\colon\mathcal{R}^{\uparrow}T^{*}M\to\mathcal{T}^{[\nabla]}.$
@@
Given a conformal connection $\nabla^{[g]}$ for $[g],$ by Proposition 3.2, the 2-nondegenerate CR structure on $\mathcal{T}^{[\nabla^{[g]}]}$ has holomorphic distribution $\widehat{\mathscr{H}}\subset\mathbb{C}\otimes T\mathcal{T}^{[\nabla^{[g]}]}$ given by $\mathrm{Ker}\{\tau^{0},\zeta^{1},\zeta^{2},\zeta^{3}\}$. Since the conformal connection $\nabla^{[g]}$ corresponds to a section $s\colon\mathcal{C}_{0}\to\mathcal{C}$, the projective connection $\nabla^{[g]}\in[\nabla^{[g]}]$ in turn defines an inclusion of principal bundles
@@
where $(\mathcal{P}\to M,\varphi)$ is the Cartan geometric data for the oriented projective structure $\nabla^{[g]}].$
@@
(4.6) $\iota(\mathcal{T}^{[g]}):=\iota(\mathcal{C}_{0})/\penalty 50\mathrm{CO}(2,\mathbb{R})\subset\mathcal{T}^{[\nabla^{[g]}]}\cong\mathcal{P}/\penalty 50P_{J},$
wherein we are using the induced injection $\mathrm{CO}(2,\mathbb{R})\to P_{J};$ see § 3.3. Recalling that $\mathcal{T}^{[g]}\cong\mathcal{R}^{\uparrow}T^{*}M,$ (4.6) defines a section $\mathcal{R}^{\uparrow}T^{*}M\to\mathcal{T}^{[\nabla^{[g]}]}.$
The pull-back of $\varphi$ for $\nabla^{[g]}]$ to $\mathcal{R}^{\uparrow}T^{*}M\subseteq\mathbb{P}_{+}T^{*}M$ is expressed as (4.2). Using the definition of the contact 1-form $\tau^{0}$ and the holomorphic 1-forms $\zeta^{i}$ from (3.13), one obtains
@@
which, by (2.17), coincides with the contact and holomorphic 1-forms for the CR structure on $\mathcal{T}^{[g]}.$ In other words, $\iota(\mathcal{T}^{[g]})$ is a Levi-nondegenerate hypersurface-type CR submanifold of $\mathcal{T}^{[\nabla^{[g]}]}$ that is transverse to its Levi kernel.
@@
A Weyl structure $\nabla^{[g]}$ for which $\mathsf{P}_{ij}=\mathsf{P}_{ji},$ is called a closed Weyl structure since, by structure equations (2.4), $\mathrm{d}\theta^{0}=0$ holds. Furthermore, given any section $s_{0}\colon M\to\mathcal{C}_{0},$ one has $\nabla^{[g]}s_{0}^{*}g=-2s_{0}^{*}\theta^{0}\otimes s_{0}^{*}g$ for $g\in\mathrm{Sym}^{2}(T^{*}\mathcal{C})$ in (2.2). Denoting $g_{0}=s_{0}^{*}g$ and $\phi_{0}=s^{*}_{0}\theta^{0},$ from $\mathrm{d}\phi_{0}=0,$ locally, $\phi_{0}=\mathrm{d}f$ holds for some function $f.$ It is then elementary to show that $\nabla^{[g]}g_{1}=0$ where $g_{1}=e^{2f}g_{0}$ and that $\nabla^{[g]}$ is the Levi-Civita connection of $g_{1}.$ In other words, closed Weyl structures are locally the Levi-Civita connection of representatives in $[g],$ defined up to positive homotheties. The Rho tensor $\mathsf{P}_{ij}$ and the Ricci tensor $R_{ij}$} of the metric are related via $\mathsf{P}_{ij}=\tfrac{-1}{n-2}({R}_{ij}-\tfrac{1}{2(n-1)}R\varepsilon_{ij})$ where $R=\varepsilon^{ij}{R}_{ij}$ and $n=\mathrm{dim}M$
@@ 4.2. CR submanifolds as holomorphic sections of the Levi kernel
To show that the conditions in Proposition 4.1 are also sufficient, we need the following definition and an observation about Weyl structures for twistor CR manifolds wherein $\mathbb{P}_{+}T^{*}M$ denotes the bundle of rays of 1-forms. We recall from § 3 that the Cartan geometry of a projective structure $\nabla$] is denoted as $(\mathcal{P}\to M,\varphi)$ has correspondence space $(\mathcal{P}\to\mathbb{P}_{+}T^{*}M,\varphi)$ equipped with an almost para-CR structure and a 7-dimensional twistor bundle $\mathcal{T}^{[\nabla]}\to\mathbb{P}_{+}T^{*}M$ equipped with a 2-nondegenerate CR structure.
@@
Given an oriented 3-dimensional projective structure $\nabla$] with twistor bundle $\mathcal{T}^{[\nabla]}\to\mathbb{P}_{+}T^{*}M,$ a section $\tilde{t}\colon N\to\mathcal{T}^{[\nabla]}$ for an open subset $N\subset\mathbb{P}_{+}T^{*}M$ is called holomorphic if the complex Pfaffian system $s^{*}\mathcal{I}=s^{*}\{\tau^{0},\zeta^{1},\zeta^{2},\zeta^{3}\},$ as defined in (3.13), is integrable, where $s=\tilde{s}\circ\tilde{t}\colon N\to\mathcal{P},$ for some section $\tilde{s}\colon\mathcal{T}^{[\nabla]}\to\mathcal{P}$.
In Remark 4.5 we will comment on how (4.12) implies that a holomorphic section, as defined above, is independent of the section $\tilde{s}\colon\mathcal{T}^{[\nabla]}\to\mathcal{P}.$ By the construction in § 3.3, the fibers of $\mathcal{T}^{[\nabla]}\to\mathbb{P}_{+}T^{*}M$ are the leaves of the Pfaffian system $\{\tau^{0},\zeta^{1},\overline{\zeta}^{1},\zeta^{2},\overline{\zeta}^{2}\}.$ It follows that, for any section $\tilde{t}\colon N\to\mathcal{T}^{[\nabla]},$ $N\subset\mathbb{P}_{+}T^{*}M,$ the 1-forms $(s^{*}\tau^{0},s^{*}\zeta^{1},s^{*}\zeta^{2},s^{*}\overline{\zeta}^{1},s^{*}\overline{\zeta}^{2})$ give a coframe on $N$ and
(4.8) $s^{*}\zeta^{3}=a_{0}\tau^{0}+a_{1}\zeta^{1}+a_{2}\zeta^{2}+b_{1}\overline{\zeta}^{1}+b_{2}\overline{\zeta}^{2},$
for some complex-valued functions $a_{0},a_{1},a_{2},b_{1},b_{2}$ on $N,$ where we have dropped $s^{*}$ on the right hand side. Since $\{\tau^{0},\zeta^{1},\zeta^{2},\zeta^{3}\}$ is integrable on $\mathbb{C}\otimes T^{*}\mathcal{T}^{[\nabla]},$ by Definition 4.3, $\tilde{t}\colon N\to\mathcal{T}^{[\nabla]}$ is holomorphic if and only if
(4.9) $b_{1}=b_{2}=0.$
@@
Recall from § 3.2 that an oriented projective structure, $\nabla$], defines an almost para-CR structure on $\mathbb{P}_{+}T^{*}M,$ which is a Cartan geometry $(\mathcal{P}\to\mathbb{P}_{+}T^{*}M,\varphi)$ of type $(\mathrm{SL}(4,\mathbb{R}),P_{1,3}),$ where $P_{1,3}=P_{0}\ltimes P_{+}$ and $P_{0}=\mathbb{R}^{+}\times\mathrm{GL}^{+}(2,\mathbb{R}).$ Since by (3.11) the fibers of the twistor bundle $\mathcal{T}^{[\nabla]}\to\mathbb{P}_{+}T^{*}M$ are $\mathrm{SL}(2,\mathbb{R})/\penalty 50\mathrm{SO}(2,\mathbb{R})\cong\mathbb{D}^{2},$ a section $\tilde{t}\colon N\to\mathcal{T}^{[\nabla]},N\subset\mathbb{P}_{+}T^{*}M,$ is equivalent to a reduction of the structure group of the almost para-CR structure on $N$ from $P_{0}=\mathbb{R}^{+}\times\mathrm{GL}^{+}(2,\mathbb{R})$ to $\widetilde{P}_{0}:=\mathbb{R}^{+}\times\mathrm{CO}(2,\mathbb{R}).$ As a result of such reduction, one obtains an inclusion $\iota_{0}\colon\widetilde{\mathcal{P}}_{0}\to\mathcal{P}$ where $\widetilde{\mathcal{P}}_{0}$ is a principal $(\widetilde{P}_{0}\ltimes P_{+})$-bundle. Moreover, by Definition 4.3 and the subsequent discussion, since $\tilde{t}\colon N\to\mathcal{T}^{[\nabla]}$ is holomorphic, it follows from (4.8) and (4.9) that
(4.10) $\tilde{t}^{*}\zeta^{3}=a_{0}\tau^{0}+a_{1}\zeta^{1}+a_{2}\zeta^{2},$
for some $a_{0},a_{1},a_{2}\in C^{\infty}(\widetilde{\mathcal{P}}_{0},\mathbb{C}).$ One can explicitly find the dependency of $a_{i}$’s on $\widetilde{P}_{0}\ltimes P_{+}$ as follows. If two points $p,q\in\widetilde{\mathcal{P}}_{0}$ lie on the fiber over $x\in N\subset\mathbb{P}_{+}T^{*}M,$ then one has $q=\mathbf{g}^{-1}p,$ for some $\widetilde{P}_{0}\ltimes P_{+}$-valued $\mathbf{g}$ given as
@@
$\varphi(p)\to\varphi(\mathbf{g}^{-1}p)=\mathbf{g}^{-1}\varphi\mathbf{g}+\mathbf{g}^{-1}\mathrm{d}\mathbf{g},$
where $\mathbf{g}\in C^{\infty}(\widetilde{\mathcal{P}}_{0},\widetilde{P}_{0}\ltimes P_{+}).$ It is straightforward to find the transformations $\tau^{0}(p)\to\tau^{0}(\mathbf{g}^{-1}p)$ and $\zeta^{i}(p)\to\zeta^{i}(\mathbf{g}^{-1}p)$ to be
[…]
@@
$\iota^{*}_{01}\zeta^{3}=a_{0}\tau^{0},$
for some $a_{0}\in C^{\infty}(\mathcal{P}_{1},\mathbb{C})$ where $\iota_{01}=\iota_{0}\circ\iota_{1}\colon\mathcal{P}_{1}\to\mathcal{P}.$ Setting $a_{1}=a_{2}=0$ in differential relations (4.13), and using the complex coframe (3.13) one arrives at
(4.14) $\iota_{01}^{*}\mu_{12}=$ $X_{0}\tau^{0}+X_{1}\zeta^{1}-Y_{1}\zeta^{2}-\textstyle{\frac{1}{2}}a_{0}\overline{\zeta}^{2}+\textstyle{\frac{1}{2}}(W^{1}_{212}+W^{2}_{121}+\mathrm{i}W^{1}_{112}+\mathrm{i}W^{2}_{221})\overline{\zeta}^{1},$
@@
for some complex-valued functions $X_{0},X_{1},Y_{0},Y_{1},Y_{2}$ on $\mathcal{P}_{1}.$ Note that to arrive at (4.14) one needs to complete the expressions (4.13) by finding the differential relations among all coframe derivatives of $a_{i}$’s and $X_{i}$’s, $Y_{i}$’s and the projective Weyl curvature. This can be done by replacing the real and imaginary parts of $\zeta^{3}$ according to (4.10) in the Cartan connection $\varphi$ of the projective structure and making sure the structure equations (3.2) remain valid. Carrying out the relevant computations, although tedious, is straightforward. It is rather remarkable that the expressions for $\mathrm{d}a_{1}$ and $\mathrm{d}a_{2},$ once pulled-back to $\mathcal{P}_{1}$ result in simple relations (4.14).
@@
Thus, one obtains a sub-bundle $\iota_{2}\colon\mathcal{P}_{2}\to\mathcal{P}_{1},$ which is a principal $\mathrm{CO}(2,\mathbb{R})\times\mathbb{R}^{+}$-bundle, defined as
(4.15) $\mathcal{P}_{2}=\left\{p\in\mathcal{P}_{1}\ \vline\ \Re Y_{1}(p)=0\right\}.$
Any holomorphic section $\tilde{t}\colon N\to\mathcal{T}^{[\nabla]},N\subset\mathbb{P}_{+}T^{*}M,$ defines a nondegenerate CR structure of hypersurface type on $N$ since $\mathscr{F}:=\mathrm{Ker}\{\tau^{0}\}$ is the canonical contact distribution and $\mathscr{H}:=\mathrm{Ker}\{\tilde{t}^{*}\mathcal{I}\}\subset\mathbb{C}\otimes\mathscr{F}$ is an integrable holomorphic distribution of a complex structure which by (3.15) satisfies the compatibility condition. In terms of the Cartan geometric description of Theorem 2.5, for the induced CR structure on $N\subset\mathbb{P}_{+}T^{*}M,$ the principal $\widetilde{P}_{0}$-bundle $\mathcal{P}_{2}\to N$ defines a CR (Weyl) connection $t\colon\mathcal{Q}_{0}\to\mathcal{Q}$ on $N$. Setting $\iota_{012}:=\iota_{0}\circ\iota_{1}\circ\iota_{2}\colon\mathcal{P}_{2}\to\mathcal{P},$ one can express the entries of $\iota_{012}^{*}\gamma$ in (2.8) as
@@
The expression for $C_{4}$ in (4.16) is given in terms of the first jet the fundamental torsion (3.7) for the almost para-CR structure on $\mathbb{P}_{+}T^{*}M$ defined by $\nabla$], which leads to Corollary 5.1. Furthermore, it follows from (4.12) that the property of being holomorphic in Definition 4.3 for a section $\tilde{t}\colon N\to\mathcal{T}^{[\nabla]},N\subset\mathbb{P}_{+}T^{*}M$ is invariant under the action of $\tilde{P}_{0}\ltimes P_{+}$, and, therefore, is independent of the section $\tilde{s}\colon\mathcal{T}^{[\nabla]}\to\mathcal{P}$. Lastly, it is clear from $C_{0}=0$ in (4.16) that the binary quartic (2.12) has a real root at $a_{1}:a_{2}]=[1:0]$ which, by the parametrization (2.11) and Proposition 2.9, corresponds to the fibers of $\mathbb{P}_{+}T^{*}M\to M.$